OpenStudy (anonymous):

the sum of 8 and a number, divided by 5 how do i write it in a phrases & simplify it? is this the phrase? or am i wrong? 8+x/5

OpenStudy (asnaseer):

almost

OpenStudy (asnaseer):

it should be (8+x)/5

OpenStudy (anonymous):

and how do i simplify that? can you show me in steps so i can understand?

OpenStudy (anonymous):

plz

OpenStudy (asnaseer):

it is already simplified, but you can expand it to this if you wish:\[\frac{8+x}{5}=\frac{8}{5}+\frac{x}{5}=1\frac{3}{5}+\frac{x}{5}\]